OpenStudy (otonogold):

Check My answer?? what is the value of log 64 8? Would it be 2 right?

OpenStudy (agent0smith):

No, you're looking for what power you'd need to raise 64 to, to get 8 64^2 will not be equal to 8. You want to get the 8, to look like 64 to the power of something (since log64 64 = 1) 8 is the square root of 64, right? So... log64 8 = log64 (64)^0.5 Can you solve it now?

OpenStudy (otonogold):

Im still kind of confused.

OpenStudy (agent0smith):

Understand this? log64 8 = log64 (64)^0.5 Since 8 is the square root of 64. It'd help if i could draw it, but OS prob won't let me.

OpenStudy (agent0smith):

The 0.5 exponent is the same as a square root, btw.

OpenStudy (otonogold):

Ohhhh okay I got it. hold on

OpenStudy (otonogold):

My answer choices: 2/3 1/2 <- my answer. 3 2

OpenStudy (agent0smith):

Yes, since log64 (64)^0.5 = 0.5 log64 64 = 0.5

OpenStudy (agent0smith):

Or think of this as log64 8 "what power do I need to raise 64 to, to get 8?" Or in general, loga b "what power do I need to raise a to, to get b?"

OpenStudy (agent0smith):

eg. log2 16 what power do you need to raise 2, to get 16? 2^5 = 16, so... 5. Or write it as log2 2^5 = 5*log2 2 = 5